The diameter of the moon is approximately one fourth the diameter of Earth, hence the volume of the moon is one fourth the volume of Earth. Is it true or false?
Answer by stanbon(48558)

The diameter of the moon is approximately one fourth the diameter of Earth, hence the volume of the moon is one fourth the volume of Earth. Is it true or false?
--------------
d(e) = 4(d(m)) where e is earth,m is moon, d is diameter
------------------
V(e) = (4/3)pi(4d(m)/2)^3 = (4/3)pi(8(d(m)^3))
V(m) = (4/3)pi(d(m)/2)^3 = (4/3)pi((1/8)(d(m))^3
V(e)/V(m) = 8/(1/8) = 64
The volume of the earth is 64 times the volume of the moon.
==============
Cheers,
Stan H.